OpenStudy (anonymous):

Damien sets up a bucket to catch raindrops. The bucket can hold 1 gallon. If each raindrop is made of about 1/90 fluid ounces of water , how many raindrops will the bucket be able to hold?

OpenStudy (anonymous):

1 US fluid ounce = 0.0078125 US gallons x fluid ounce = 1 gallon. Solve for x. x = 1/0.0078125 = 128. Then: 1/90 in a drop 128 in y drops, solve for y. y = 128*90 = 11520 rain drops. If I did not make any careless mistake. :-)

OpenStudy (anonymous):

thank you :)